OpenStudy (anonymous):

If f(x)= x/x-3 and g(x)= 2x -1, find (f-g)(x)

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{x}{x-3}\qquad g(x)= 2x -1 \\ \quad \\ (f-g)(x)\implies f(x)-g(x)\implies \cfrac{x}{x-3}\quad -\quad 2x -1\)

OpenStudy (anonymous):

and then what? i got that far but the ending result needs to be a fraction @jdoe0001

OpenStudy (jdoe0001):

yes.... well.. get the LCD and add them up

OpenStudy (jdoe0001):

\(\bf f(x)-g(x)\implies \cfrac{x}{x-3}\quad -\quad 2x -1\implies \cfrac{x}{x-3}\quad -\quad \cfrac{2x}{1}-\cfrac{1}{1}\) got LCD?

OpenStudy (anonymous):

ugh no. is it x-3 or is it 1

OpenStudy (anonymous):

or none of those

OpenStudy (jdoe0001):

hmmm well.... is a "value you can divide all those by" so in short , if not found quickly you can always just multiply them all and and that'd be the LCD so \(\bf (x-3)\cdot 1\cdot 1\implies 1(x-3)\implies (x-3)\)

OpenStudy (anonymous):

i got it thanks! :)

OpenStudy (jdoe0001):

yw